A sample of chlorine gas occupies a volume of 946mL at a pressure 726mmHg. Calculate the pressure of the gas if the volume is reduced to a quarter at constant temperature


1
Expert's answer
2021-01-28T05:38:19-0500

find the formula:

"P1V1=P2V2" at constant temperature


"P2=\frac{P1V1}{V2}"


"V2=\frac{V1}{4}=\frac{946}{4}=236.5"


"P2=\frac{\frac{726}{760}\times 946}{236.5}=3.82" atm